# Gaming on Coincident Peak Shaving: Equilibrium and Strategic Behavior

### v4

@@ VIII Discussion and Conclusion
Note that the best response of both the virtual CP agent and CP-period agent set is to shift demand away from the CP period, and the benefits they obtained monotonically increase with the shifting amount. From (72), we know the maximum shifting amount of the CP-period agent set is less than the best strategy of the virtual CP agent, indicating the best strategy for the CP-period agent set is their maximum shifting amount $\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}$. In this case, the best strategy of agents in the CP-period agent set is their maximum demand shifting, i.e., $x_{i}^{*}=\min\{r_{i},b_{i}\},i\in N_{\mathrm{cp}}$
We then verify in this case the non-CP-period agent set will still balance system demand to get profits. We first show the non-concave game condition still holds, which is equivalent between virtual agents and agent set, i.e.,
$-\sum_{i\in N_{\mathrm{ncp}}}r_{i}-\sum_{i\in N_{\mathrm{cp}}}r_{i}=-\sum_{i\in N}r_{i}\leq b\leq\sum_{i\in N_{\mathrm{ncp}}}r_{i}+\sum_{i\in N_{\mathrm{cp}}}r_{i}=\sum_{i\in N}r_{i}$ (73)
Thus, under the condition (72) with $\sum_{i\in N_{\mathrm{cp}}}x_{i}^{*}=\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}$, the non-CP-period agent set can still balance system demand. Then, we show the best response rationale is the same between the non-CP-period agent set and the virtual non-CP agent due to the same baseline demand and critical point conditions. From Theorem 5, we know the virtual non-CP agent benefited by shifting demand to the CP period in response to the virtual CP agent’s strategy. Thus, the benefit of the non-CP-period agent set increases with the shifting amount to the CP period because it offsets the shifting of the CP-period agent set and further fills the system demand difference when the CP-period-agent set reaches maximum shifting capability. Also, due to $\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}<\sum_{i\in N_{\mathrm{cp}}}r_{i}$, we have $b-\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}>b-\sum_{i\in N_{\mathrm{cp}}}r_{i}$, meaning that the non-CP-period agent set can actually get more benefit than the virtual non-CP agent.
To conclude, the CP-period agents in the CP-period agent set with strategy $x_{i}^{*}=\min\{r_{i},b_{i}\},i\in N_{\mathrm{cp}}$ form a best response to the aggregated response of non-CP-period agent set with strategy $\sum_{i\in N_{\mathrm{ncp}}}x_{i}^{*}=b-\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}$ under the condition of
$\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}<\sum_{i\in N_{\mathrm{cp}}}b_{i},-\sum_{i\in N}r_{i}\leq b\leq\sum_{i\in N}r_{i};$ (74)
@@
$\sum_{i\in N_{\mathrm{ncp}}}\max\{-r_{i},b_{i}\}>\sum_{i\in N_{\mathrm{ncp}}}b_{i},-\sum_{i\in N}r_{i}\leq b\leq\sum_{i\in N}r_{i}.$ (76)
Note that here, the set of agents is fixed, and we focus on the set level performance for the CP-period agent set and non-CP-period agent set. Specifically, when the aggregated (set-level) demand of the CP-period agent set is higher in the non-CP period, the CP-period agent set will perform as the non-CP-period agent set, and meanwhile, the aggregated demand of the non-CP-period agent set must be higher in the CP period, and the non-CP-period agent set perform as the CP-period agent set. This is similar to the two-agent case, where they only exchange strategy when the relative relationship between their own peak period and CP period changes.

### v5

@@ VIII Discussion and Conclusion
Since both the virtual CP agent and the CP-period agent set benefit from shifting demand away from the CP period, their best response is to maximize their shifting amount, as their cost reduction monotonically increases with demand shifting. From (68), we observe that the total maximum shifting capacity of the CP-period agent set is lower than the optimal shifting amount of the virtual CP agent. This implies that the best strategy for the CP-period agent set is to fully utilize its shifting capacity, leading to an equilibrium shifting amount of $\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}$. Consequently, each agent within the CP-period agent set follows its individual maximum shifting strategy, resulting in $x_{i}^{*}=\min\{r_{i},b_{i}\},i\in N_{\mathrm{cp}}$.
We then verify that, in this case, the non-CP-period agent set will still balance system demand to maximize its profits. First, we confirm that the non-concave game condition remains valid. Since the conditions for the virtual agents and the agent sets are equivalent, i.e.,
$0\leq b\leq\sum_{i\in N_{\mathrm{ncp}}}r_{i}+\sum_{i\in N_{\mathrm{cp}}}r_{i}=\sum_{i\in N}r_{i}$ (69)
Thus, under the condition (68), where $\sum_{i\in N_{\mathrm{cp}}}x_{i}^{*}=\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}$, the non-CP-period agent set can still balance system demand. Next, we show that the best response rationale remains the same between the non-CP-period agent set and the virtual non-CP agent due to their identical baseline demand and critical point conditions. From Theorem 1, we know that the virtual non-CP agent benefits by shifting demand to the CP period in response to the virtual CP agent’s strategy. Similarly, the benefit of the non-CP-period agent set increases as it shifts more demand to the CP period. This not only counteracts the demand shifting of the CP-period agent set but also helps to further balance system demand once the CP-period agent set reaches its maximum shifting capacity. Additionally, since $\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}<\sum_{i\in N_{\mathrm{cp}}}r_{i}$, we have $b-\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}>b-\sum_{i\in N_{\mathrm{cp}}}r_{i}$, indicating that the non-CP-period agent set can actually achieve greater benefits than the virtual non-CP agent.
To conclude, the CP-period agents in the CP-period agent set with strategy $x_{i}^{*}=\min\{r_{i},b_{i}\},i\in N_{\mathrm{cp}}$ form the best response to the aggregated response of non-CP-period agent set with strategy $\sum_{i\in N_{\mathrm{ncp}}}x_{i}^{*}=b-\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}$ under the condition of
$\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}<\sum_{i\in N_{\mathrm{cp}}}b_{i},0\leq b\leq\sum_{i\in N}r_{i};$ (70)
@@
$\sum_{i\in N_{\mathrm{ncp}}}\max\{-r_{i},b_{i}\}>\sum_{i\in N_{\mathrm{ncp}}}b_{i},0\leq b\leq\sum_{i\in N}r_{i}.$ (72)
Note that here, the set of agents remains fixed, and we focus on the aggregate (set-level) performance of both the CP-period agent set and the non-CP-period agent set. Specifically, when the total demand of the CP-period agent set is higher in the non-CP period, this set effectively behaves like the non-CP-period agent set. Simultaneously, the total demand of the non-CP-period agent set must be higher in the CP period, causing it to act as the CP-period agent set. This dynamic mirrors the two-agent case, where agents only swap strategies when the relative relationship between their individual peak periods and the system CP period changes.